OpenStudy (anonymous):

Identify the vertex for the graph of y = 2x^2 + 8x − 3

hero (hero):

x = -b/2a = -8/2(2) = -8/4 = -2 2(-2)^2 + 8(-2) - 3 = 8 - 16 - 3 = 8 - 19 = -11 (2,-11)

hero (hero):

(-2,-11)

hero (hero):

Mandolino

OpenStudy (anonymous):

the form y=a(x-h)^2+k has vertex (h,k) and can be found by completing the square\[y=2(x^2+4x+4)-3-8\]\[y=2(x+2)^2-11\]vertex is (-2, -11)

hero (hero):

Why on earth would you do all of that work?

OpenStudy (anonymous):

i didn't know i was in a race with hero; i thought you left

OpenStudy (anonymous):

also, i find the other way tedious

hero (hero):

In what way?

hero (hero):

It's by far the easiest way

OpenStudy (anonymous):

all that squaring and multiplying and adding and subtracting... it way too much like maths

hero (hero):

Oh, I get it. It's not "systematic" or "algebraic". Kind of reminds me of why I hate the quad formula

OpenStudy (anonymous):

here we go again...

hero (hero):

And your way is better in what way? You do the same things...

OpenStudy (anonymous):

just a preference; we better be careful we may get called the orifice of a digestive system again

OpenStudy (anonymous):

the end, not the beginning of the system

hero (hero):

WTH Mandolino

OpenStudy (anonymous):

don't you remember that guy yesterday when i wrote kids these days

hero (hero):

I don't remember being called anything. I must have overlooked that part

OpenStudy (anonymous):

Determine whether the graph of y = 4x2 + 5x − 8 opens up or down and whether it has a maximum or minimum point. Up; Minimum Up; Maximum Down; Minimum Down; Maximum

OpenStudy (anonymous):

when the lead coefficient, 4 in this problem, is positive the parabola opens upward and thus the vertex is a minimum

OpenStudy (anonymous):

Identify the solutions of 3x^2 + 14x + 8 = 0.

OpenStudy (anonymous):

(3x+2)(x+4)=0 x=-2/3 or x=-4

OpenStudy (anonymous):

Solve x^2 − 7x + 12 = 0

OpenStudy (anonymous):

(x-4)(x-3)=0 x=4 or x=3

hero (hero):

That's not systematic Mandolino

OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 30 = x^2 + x? (−5, 0) and (6, 0) (−5, 0) and (−6, 0) (5, 0) and (−6, 0) (5, 0) and (6, 0)

hero (hero):

Safari King must be one of those online testers

OpenStudy (anonymous):

y=x^2+x-30=(x+6)(x-5)=0 x=-6 or x=5

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

a lot of HW

OpenStudy (anonymous):

Just a person in need of MAJOR MATH HELP

OpenStudy (anonymous):

Where does the graph of y = 3x^2 − 10x − 8 cross the x-axis?

OpenStudy (anonymous):

any others, I've got about 7 minutes

OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand of the quadratic formula is a positive number that is not a perfect square? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

What are the exact solutions of x^2 − 5x − 7 = 0?

OpenStudy (anonymous):

y=3x^2-10x-8=(x-4)(3x+2)=0 x=4 or x=-2/3 the points are (4,0); (-2/3,0)

OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand of the quadratic formula is a positive number that is not a perfect square? answer = two irrational solutions

OpenStudy (anonymous):

What are the approximate solutions of 2x^2 + x = 14 rounded to the nearest hundredth? x ≈ −2.41 and x ≈ 2.91 x ≈ −2.91 and x ≈ 2.41 x ≈ −5.82 and x ≈ 4.82 No real solutions

OpenStudy (anonymous):

you have to use quadratic formula to solve What are the exact solutions of x^2 − 5x − 7 = 0?

OpenStudy (anonymous):

Just keep that question of exact solutions

OpenStudy (anonymous):

skip*

OpenStudy (anonymous):

discriminat=25-4(-7)=53 then\[x=\frac{5 \pm \sqrt{53}}{2}\]

OpenStudy (anonymous):

Solve 3x2 + 5x + 7 = 0. Round solutions to the nearest hundredth. No real solutions x ≈ −6.34 and x ≈ 1.34 x ≈ −0.45 and x ≈ 2.11 x ≈ −2.11 and x ≈ 0.45

OpenStudy (anonymous):

What are the approximate solutions of 2x2 + x = 14 rounded to the nearest hundredth? x ≈ −2.41 and x ≈ 2.91 x ≈ −2.91 and x ≈ 2.41 x ≈ −5.82 and x ≈ 4.82 No real solutions

OpenStudy (anonymous):

What are the approximate solutions of 2x^2 + x = 14 rounded to the nearest hundredth? 2x^2+x-14=0 D=1-4(2)(-14)=113 so\[x=\frac{-1 \pm \sqrt{113}}{4}\] Answer is x ≈ −2.91 and x ≈ 2.41

OpenStudy (anonymous):

Solve 3x2 + 5x + 7 = 0. Round solutions to the nearest hundredth. D=25-4(2)(7)<0 so the answer is No real solutions

OpenStudy (anonymous):

A firecracker shoots up from a hill 145 feet high with an initial speed of 80 feet per second. Using the formula H(t) = −16t2 + vt + s, approximately how long will it take the firecracker to hit the ground?

OpenStudy (anonymous):

i think i got them all

OpenStudy (anonymous):

Darlene kicks a soccer ball off the ground and in the air with an initial velocity of 34 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches?

hero (hero):

Are you going to make us take the whole test?

OpenStudy (anonymous):

i think you will have to repost; i only had until the top of the hour; sorry

OpenStudy (anonymous):

Using graphing technology, approximate the solutions of 0.35x2 − 2.52x + 1.34 = 0. x ≈ 0.12 and x ≈ 4.56 x ≈ 0.58 and x ≈ 6.62 x ≈ 1.17 and x ≈ 6.14 x ≈ 1.83 and x ≈ 4.84

OpenStudy (anonymous):

its just HW i'm pretty sure

OpenStudy (anonymous):

Oh okay. Well Thanks Anyways You were a HUGE help

OpenStudy (anonymous):

later SK and Hero

hero (hero):

Mandolino, you're too nice and too assuming

OpenStudy (anonymous):

:})

OpenStudy (anonymous):

Yeah he's something you need to be hero

hero (hero):

:P

OpenStudy (anonymous):

gn

hero (hero):

I'm already a hero

hero (hero):

In fact, I'm a master :P

OpenStudy (anonymous):

and a master